Calculate the pH and pOH of 1.2 x 10-3 M HCl solution. Calculate pH, pOH and...

  1. Calculate the pH and pOH of 1.2 x 10-3 M HCl solution.
  1. Calculate pH, pOH and [OH-] of 0.1 M HNO3 solution.

  1. If a solution X has pH = 5, which of the following is true:
  1. Solution X is neutral.
  2. H3O+ ion concentration is higher than OH- concentration.

c. OH- ion concentration is higher than H3O+ concentration.
Answer #1

Hell: 1.2x163M pH- - log[H + 1 : -lo 1.2*1003) pH = 2.92 POH= 14-04 . 14-2.92 POH : 11.08 HNO3 PH= -log (6.1) pona 14-1 = 13